149
edits
No edit summary |
No edit summary |
||
--[[User:Zjsimmons|Zjsimmons]] ([[User talk:Zjsimmons|talk]])
Week in review: I've basically been working on 2 things, trying to get better lock performance and starting to write up the EIT eqns so we can look at that in the context of the green transition in Eu:YSO. Nick and I tried to look for absorption on thur but didn't see very good results; stability of the output power/lock issues are really fighting us. We did learn some things though:
*tuning up alignment to get even photodiode response seemed to help the lock performance, perhaps it helps with noise to have the optical feedback balanced?
*the quality of the laser makes a big difference to the lock performance. i.e. the laser may look single mode on the spectrum analyzer and the wavemeter reading may be very stable but touching the current or feedback makes the error signal look much better. When it's poor it may appear that a band of noise periodically passes through the peaks, disturbing everything. zooming in on the peak, it looks noisy.
Today i made some changes to the electronics.
* '''Added 50 ohm impedance matching resistors to the photodiode inputs.''' I had been hesitant to do this on account of it shrinking our input signal but decided to try it on account of noise; perhaps the low impedance would be less susceptible from noise coming from the photodiodes or noise from some ground loop through the photodiodes. Indeed the error signal shrank somewhat, but was still plenty big enough to do feedback.
* '''Added a new passive fast feedback branch that goes directly to the laser diode.''' I talked to Larry yesterday about our issues trying to feedback through the laser driver. He said that that's problematic for same two reasons we had been suspecting: ground/noise issues and the fact the laser driver mod input has very high gain, so you need very low noise. He suggested a fast feedback branch directly to the diode and sketched out a little circuit. I implemented something like it. Turns out we had a lot of little discarded fast feedback boxes lying around so i cannibalized one of those used the layout suggested by Larry.
--[[User:Zjsimmons|Zjsimmons]] ([[User talk:Zjsimmons|talk]]) 18:33
Made some changes today, see qualitative improvement in green lock. So we've been having issues with the locking performance, the transmitted IR and also generated green were intensity modulated at about 300Hz, a strange freq to see problems as it was not the cavity piezo resonance freq of about 420Hz. The SHG derived green looked at times almost like a pulse train as a result of the modulation. This was tracked to in part a suspected ground loop between driver and locking electronics. Unfortunately the fast feedback was needed to get good output power, but would also cause the modulation.
*Swapped the cavity piezo for the new shorter travel (6um) one from Noliac. Glued everything together with epoxy so as to cut down on load weight. Somehow, realigning the cavity didn't turn into a huge project. Note: when you see some green, put the spots on top of each other and you're most of the way toward good alignment.
|